# SUMO: Subspace-Aware Moment-Orthogonalization for Accelerating Memory-Efficient LLM Training

### v1

@@ Acknowledgment
TT was supported by the Israel Science Foundation (No. 1940/23) and MOST (No. 0007091) grants.
@@ Appendix A Proofs of Section 3
Lemma A.1 (Moment Becomes Low-Rank During Training).
Let ${\bf M}^{(t)}\in\mathbb{R}^{n\times m}$ denote the first moment of a reversible layer in a moment-based optimization algorithm, updated according to
${\bf M}^{(t)}=\beta{\bf M}^{(t-1)}+{\bf G}^{(t)},$
where ${\bf G}^{(t)}$ is the gradient matrix at iteration $t$. Let ${\bf M}^{(t)}={\bf U}^{(t)}\boldsymbol{\Sigma}^{(t)}{{\bf V}^{(t)}}^{\top}$ be the singular value decomposition (SVD) of ${\bf M}^{(t)}$, and define the rank-$r$ orthogonal projection matrix as ${\bf P}^{(t)}(r)={\bf U}^{(t)}[:,1\!:\!r]\,{\bf U}^{(t)}[:,1\!:\!r]^{\top}$. Then the relative error of the best rank-one approximation,
satisfies $\kappa_{M}(t)\leq O(C^{-t})$ for some constant $C>1$.
Proof of Lemma A.1.
@@
for some constant $C>1$Box
Before proving Lemma LABEL:lma:moun_convergence, we shortly present the following two preliminary lemmas. To that end, we present the following notations,
  • ${\bf M}^{(t)}$ The moment in iteration $t$. Its dimensions are $n\times m$, where $n<m$.
[…]
  • $\mathcal{L}^{*}$ A stationary point to which the loss $\mathcal{L}$ converges.
  • $B$ Batch size.
@@
[…]
@@
Where in ()(*)( ∗ ) we used [53], equation 2.8. $\Box$
@@
Lemma 3.3 (Exact convergence rate of Muon). Consider the Muon optimizer update defined by
$\mathbf{M}^{(t)}\leftarrow\beta\mathbf{M}^{(t-1)}+(1-\beta)\mathbf{G}_{t}$
[…]
$\mathbf{W}^{(t+1)}\leftarrow\mathbf{W}^{(t)}-\eta_{t}\mathbf{O}^{(t)},$
where $\mathbf{M}^{(t)}=\mathbf{U}^{(t)}\mathbf{S}^{(t)}\mathbf{V}^{(t)^{\top}}$ denotes the singular value decomposition of $\mathbf{M}^{(t)}$, and $\mathcal{E}_{i}^{(t)}$ represents the Newton-Schulz5 approximation error after $i$ iterations. Assuming A(1) - A(4), If we take $\beta=1-\alpha$ with $\alpha=\min(\frac{\sqrt{Rl}}{\sigma\sqrt{T}},1)$, $\eta_{t}=\eta=\frac{\sqrt{4R}}{\sqrt{(10/(1-\beta)+2m+4m\delta+2m\delta^{2})TL}}$, and $B=1$ (batch free convergence) than $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\|\nabla\mathcal{L}(\mathbf{W}^{(t)})\|\right]$ is bounded by
where $R=\mathcal{L}(\mathbf{W^{(0)}})-\mathcal{L}^{*}$. If we take $\beta$ as an arbitrary constant, we have to take $B=T$, and we have,
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla\mathcal{L}(\mathbf{W}^{(t)})\|\leq\mathcal{O}\left(\left[\frac{\sqrt{RLn(2+4\delta+2\delta^{2})}}{\sqrt{T}}+\frac{\sqrt{RL}}{\sqrt{T}}+\frac{\sigma}{T^{3/2}}+\frac{\sigma}{\sqrt{T}}\right]\frac{1}{1-4\sqrt{m}\delta}\right).$
Proof of Lemma Lemma LABEL:lma:moun_convergence.
@@
Lemma 3.2 (Orthogonalization error $\mathbf{\mathcal{E}}_{i}$) For a matrix ${\bf A}\in\mathbb{R}^{m\times n}$, let $\sigma_{1}$ be the largest singular value of ${\bf A}{\bf A}^{\top}$ and $\sigma_{m}$ be the smallest (without the loss of generality, assume $m\leq n$). Let $r\leq m$ be the largest index where $\sigma_{r}>\sigma_{r+1}=\dots\ \sigma_{m}\geq 0$. Let $\kappa=\frac{\sigma_{1}}{\sigma_{m}}$ by the condition number of ${\bf A}{\bf A}^{\top}$. Denote $\mathbf{\mathcal{E}}_{i}$ the error of Newton-Schultz after $i$ iterations. Then we have
@@
Theorem 3.8 (Convergence of SUMO) For a loss function $\mathcal{L},$ and given architecture $\Phi$, suppose that the compositions of $f\equiv\mathcal{L}\left(\Phi(\cdot)\right)$ is $\beta$-smooth non-convex function that is bounded by some $M\in\mathbb{R}_{+}$. Let ${\bf G}^{(t)}_{j}$ denote the gradient matrix w.r.t. the $j$-th reversible layer ${\bf W}^{(t)}_{j},$ at time $t\in\mathbb{N}$, for all $j\in[L]$ and $t\in\mathbb{N}$, and $T_{\ell},\ell\in\mathbb{N}$ times are set by a convergence criterion (that is, $\|\hat{{\bf G}}^{\left(\mathsf{T}_{\ell}\right)}_{j}\|\leq\varsigma_{\ell}$). Then, there exist $\mathsf{C}\in\mathbb{R}_{+}$ and $N$ such that for all $\mathsf{T}_{N}>\frac{\mathsf{C}}{\varepsilon^{2}}$, and $\frac{1}{\mathsf{T}_{N}}\sum_{i=0}^{N-1}\sum_{t=\mathsf{T}_{i}}^{\mathsf{T}_{i+1}-1}\left\|{{\bf G}^{(t)}_{j}}\right\|_{F}^{2}\leq\varepsilon$. Namely, Algorithm 1 achieves an $\varepsilon$-critical point, [^3] Also known as $\varepsilon$-stationary, see, e.g., [14]. i.e., $\left\|{\bf G}^{(t)}_{j}\right\|_{F}^{2}\leq\varepsilon$, for some $t\in\mathbb{N}$, and any $j\in[L]$.
@@
for any layer $j\in[L]$; in the following, for simplicity of notation, we ignore the index $j$ and use ${\bf G}^{(t)}$ instead. By Lemma A.4, the low-rank optimization block 1 in Algorithm 1 is guaranteed to converge; we denote by $\mathsf{T}_{\ell}\in\mathbb{N}$ the time index $t$ at which we enter block 1 for the $\ell$th time (i.e., $\|\hat{{\bf G}}_{j}^{\left(\mathsf{T}_{\ell}\right)}\|\leq\varsigma_{2}$), for $\ell\in\mathbb{N}$. Furthermore, we recall that ${\bf G}_{j}^{(t)}\triangleq\nabla_{\boldsymbol{{\bf W}_{j}}}f\left(\boldsymbol{\theta}^{(t)}\right)$; when clear from the context, we omit $j$ from ${{\bf W}_{j}}$, and use instead $\nabla_{\boldsymbol{{\bf W}_{j}}}f\left(\boldsymbol{\theta}^{(t)}\right)=\nabla f\left({\bf W}^{(t)}\right)$. Consider the SVD decomposition of the gradient $\nabla_{\boldsymbol{{\bf W}^{j}}}f\left(\boldsymbol{\theta}_{\mathsf{T}_{i}}\right)={\bf U}^{(\mathsf{T}_{i})}\boldsymbol{\Sigma}^{(\mathsf{T}_{i})}{{\bf V}^{(\mathsf{T}_{i})}}^{\top}$. For $t\in[\mathsf{T}_{i},\mathsf{T}_{i+1}-1]$, we define the projected gradient as $\hat{{\bf G}}^{(t)}\triangleq{\bf P}^{(\mathsf{T}_{i})}(r){\bf G}^{(t)},$ where ${\bf P}^{(\mathsf{T}_{i})}(r)={\bf U}^{(\mathsf{T}_{i})}\left[:,:r\right]^{\top}$, using the exact truncated-SVD calculation (in Block 1). For simplicity, we refer to ${{\bf Q}}^{(\mathsf{T}_{i})}(r)$ as ${{\bf Q}}^{(\mathsf{T}_{i})},$ and we denote ${\bf P}^{(\mathsf{T}_{i})}={{{\bf Q}}^{(\mathsf{T}_{i})}}^{\top}{{\bf Q}}^{(\mathsf{T}_{i})}$. Next, let $h_{t}\triangleq f({\bf W}^{(t)})-f({\bf W}^{(\mathsf{T}_{i+1})})$, and $\eta_{t}$ denote the learning rate. Then,
@@
Assume a constant learning rate $\eta_{t}=\eta$, and define $T_{i}:=\mathsf{T}_{i+1}-\mathsf{T}_{i}$. For each interval $[\mathsf{T}_{i},\mathsf{T}_{i+1}-1]$, we have,
@@
Consider the same setting and assumptions as in Theorem 3.8. Then, the second time $t=\mathsf{T}_{\ell}\in\mathbb{N}$ in which Algorithm 1 enters Block 1 (where it updates the projection matrix) happens for a finite $\ell\in\mathbb{N}$
@@
$-\langle\hat{{\bf G}}^{(t)},{\bf O}^{(t)}\rangle\leq-\frac{1}{4}\|\hat{{\bf G}}^{(t)}\|_{F}^{2}+\frac{5}{2}\|\hat{{\bf G}}^{(t)}-{\bf M}^{(t)}\|_{F}^{2}.$
@@ Appendix C Update Step Rule Formulation
$\textbf{SUMO}\left\{\begin{aligned} &\hat{{\bf G}}^{(t)}={{{\bf Q}}^{(t)}}^{\top}\nabla_{\bf W}f\left({\bf W}_{t};\xi_{t}\right){{{\bf R}}^{(t)}}\quad\\ &{\bf M}^{(t+1)}=\beta{\bf M}^{(t)}+(1-\beta)\hat{{\bf G}}^{(t)}\\ &{\bf O}^{(t+1)}=\text{Orthogonalization\_SVD}\left({\bf M}^{(t+1)}\right)\\ &{\bf W}^{(T)}={\bf W}^{(0)}+\eta\sum_{t=0}^{T-1}{{{\bf Q}}^{(t)}}{\bf O}^{(t+1)}{{{\bf R}}^{(t)}}^{\top},\end{aligned}\right.$
@@ Appendix D Additional Experiments
D.1 Details of Fine-Tuning on GLUE
We fine-tune the pre-trained RoBERTa-Base model on the GLUE benchmark using the model provided by the Hugging Face. In Table and, we detail the hyper parameters used in fine-tuning.
@@ D.1 Details of Fine-Tuning on GLUE
Rank Config. $r=4$
Projection back scale 4
Max Seq. Len. 512
@@
Rank Config. $r=8$
Projection back scale 2
Max Seq. Len. 512

### v2

@@ Acknowledgment
TT was supported by the Israel Science Foundation (No. 1940/23) and MOST (No. 0007091) grants. OL was supported by the MOST grant No. 0007341.
@@ Appendix A Proofs of Section 3
Proof of Lemma 3.1.
@@
for some constant $C>1$ . $\Box$
Before proving Lemma 3.3, we shortly present the following two preliminary lemmas. To that end, we present the following notations,
  • ${\bf M}^{(t)}$ - The moment in iteration $t$. Its dimensions are $n\times m$, where $n<m$.
  • $\|\cdot\|$ - The Frobenius norm: $\|{\bf A}\|=\|{\bf A}\|_{F}=\sqrt{{\bf A}{\bf A}^{\top}}$
  • $\mathcal{L}^{*}$ - A stationary point to which the loss $\mathcal{L}$ converges.
  • $B$ - Batch size.
@@
${\bf O}^{(t)}\leftarrow{\bf U}^{(t)}{{\bf V}^{(t)}}^{\top}+\mathbf{\mathcal{E}}^{(t)}_{5},\quad\text{(Newton-Schulz 5 iteration approximation)},$
@@
Where in ()(*) we used [53], equation 2.8. $\Box$
@@
Proof of Lemma 3.3.
@@
[…]
@@
Assume a constant learning rate $\eta_{t}=\eta$, and define $T_{i}:=\mathsf{T}_{i+1}-\mathsf{T}_{i}$. For each interval $\mathsf{T}_{i},\mathsf{T}_{i+1}-1]$, we have,
@@
Consider the same setting and assumptions as in Theorem 3.8. Then, the second time $t=\mathsf{T}_{\ell}\in\mathbb{N}$ in which Algorithm 1 enters Block 1 (where it updates the projection matrix) happens for a finite $\ell\in\mathbb{N}$ .
@@
$\langle\hat{{\bf G}}^{(t)},{\bf O}^{(t)}\rangle\leq-\frac{1}{4}\|\hat{{\bf G}}^{(t)}\|_{F}^{2}+\frac{5}{2}\|\hat{{\bf G}}^{(t)}-{\bf M}^{(t)}\|_{F}^{2}.$
@@ Appendix C Update Step Rule Formulation
$\textbf{SUMO}\left\{\begin{aligned} &\hat{{\bf G}}^{(t)}={{{\bf Q}}^{(t)}}^{\top}\nabla_{\bf W}f\left({\bf W}_{t};\xi_{t}\right){{{\bf R}}^{(t)}}\qquad\\ &{\bf M}^{(t+1)}=\beta{\bf M}^{(t)}+(1-\beta)\hat{{\bf G}}^{(t)}\\ &{\bf O}^{(t+1)}=\text{Orthogonalization\_SVD}\left({\bf M}^{(t+1)}\right)\\ &{\bf W}^{(T)}={\bf W}^{(0)}+\eta\sum_{t=0}^{T-1}\Bigl({\bf G}^{(t)}-{{\bf Q}}^{(t)}\bigl(\hat{{\bf G}}^{(t)}-{\bf O}^{(t+1)}\bigr){{{\bf R}}^{(t)}}^{\top}\Bigr),\end{aligned}\right$.
@@ Appendix D Additional Experiments
Comparison with Muon.
Table 8 below complements Table 2 by comparing full fine-tuning, vanilla Muon, and our SUMO; results are reported as mean ± standard deviation when available.
Table 8: Additional comparison to Table 2: full fine-tuning vs. vanilla Muon and our SUMO (mean $\pm$ std when reported).
Model Memory CoLA STS-B MRPC RTE SST-2
Full Fine-Tuning 747M 62.24 90.92 91.30 79.42 94.57
Muon Full Fine Tuning 458M 61.19 90.98 92.14 80.83 94.71
SUMO (Newton-Schulz5, rank=4) 197M 61.81 $\pm$ 0.02 90.81 $\pm$ 0.013 92.43 $\pm$ 0.034 79.33 $\pm$ 0.031 94.14 $\pm$ 0.028
SUMO (SVD, rank=4) 197M 62.32 $\pm$ 0.015 91.05 $\pm$ 0.007 93.48 $\pm$ 0.022 81.08 $\pm$ 0.019 94.93 $\pm$ 0.01
SUMO (Newton-Schulz5, rank=8) 198M 61.73 $\pm$ 0.021 90.77 $\pm$ 0.032 91.93 $\pm$ 0.04 79.66 $\pm$ 0.03 94.13 $\pm$ 0.025
SUMO (SVD, rank=8) 198M 61.69 $\pm$ 0.014 91.11 $\pm$ 0.02 93.72 $\pm$ 0.018 81.38 $\pm$ 0.011 94.83 $\pm$ 0.01
These results show that our SUMO, achieves better performance with a significantly smaller memory footprint compared to Muon full fine-tuning approach.
Hyperparameters Grid search.
To evaluate the impact of Subspace Update Frequency (K) and Ranks (r), we performed a grid search during the pretraining of the LLaMA 130M model on the C4 dataset. This specific setup allows for a direct comparison with the Galore method. Table: Perplexity results from a grid-search of Subspace Update Frequency (K) and Ranks (r) for the LLaMA 130M model pretrained on the C4 dataset. Values are presented as Galore/SUMO.
Table 9: Perplexity from a grid-search over Subspace Update Frequency ($K$) and Rank ($r$) for LLaMA-130M on C4. Values are Galore/SUMO
Update Frequency Rank = 128 Rank = 256 Rank = 512
100 29.7/28.27 27.9/26.74 27.4/26.73
250 28.1/27.86 26.5/24.87 26.2/24.82
500 27.2/25.91 25.6/24.98 25.3/24.31
1k 26.8/25.83 25.1/25.42 24.8/24.93
D.1 Details of benchmarks in Table 6
Specifically, the pretrained models is evaluated on the following tasks:
  • Perplexity: Measured on the C4 dataset [46].
  • Commonsense Reasoning: BoolQ [56], RTE [57], HellaSwag (HS) [58], Winogrande (WG) [59], OpenBookQA (OBQA) [60], ARC-Easy (ARC-E), and ARC-Challenge (ARC-C) [61].
  • Physical and Scientific Reasoning: PIQA [62], SciQ [63], and MathQA [64].
D.2 Details of Fine-Tuning on GLUE
We fine-tune the pre-trained RoBERTa-Base model on the GLUE benchmark using the model provided by the Hugging Face. In Table 11 and, we detail the hyper parameters used in fine-tuning.
@@ D.2 Details of Fine-Tuning on GLUE
Rank Config. $r=4$ $r=4$ $r=4$ $r=4$ $r=4$ $r=4$ $r=4$ $r=4$
Projection back scale 4 4 4 4 4 4 4 4
Max Seq. Len. 512 512 512 512 512 512 512 512
@@
Rank Config. $r=8$ $r=8$ $r=8$ $r=8$ $r=8$ $r=8$ $r=8$ $r=8$
Projection back scale 2 2 2 2 2 2 2 2
Max Seq. Len. 512 512 512 512 512 512 512 512